OpenStudy (anonymous):

Differentiate the function. f(x) = log10(x^7 + 4)

OpenStudy (anonymous):

does it help to know that \[\log_{10}(x^7+4)=\frac{\ln(x^7+4)}{\ln(10)}\]?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

it should since \[\frac{d}{dx}\ln(f(x))=\frac{f'(x)}{f(x)}\] and \[\ln(10)\] is just a number

OpenStudy (anonymous):

(7x^6)/(x^7 +4)log(10) ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well not log(10) but rather \(\ln(10)\) unless that is what you meant

OpenStudy (anonymous):

yup got it thanks!!

OpenStudy (anonymous):

one more question

OpenStudy (anonymous):

@satellite73